# Diffusive contact between randomly driven colloidal suspensions

### v1

@@ Abstract
Our study highlights two key findings in this regime. First, the specific mechanism driving the effective temperature does not influence the relaxation of connected systems toward a steady state. Second, the effective temperature serves a dual purpose: it defines a generalized chemical potential, and together, these two parameters govern the thermalization process in connected systems. In other words, the steady state of the connected systems is reached at maximum entropy, as in thermal equilibrium. Remarkably, this occurs notwithstanding the continuous investment of work and the concomitant entropy production required to sustain a driven, steady state.
In this study, our focus was on driven dilute colloidal suspensions. We anticipate that our effective chemical potential may prove inadequate in describing the equilibration of diffusive contacts in scenarios where the GFDR is violated. For instance, this could occur in denser driven/active colloidal suspensions [1, 13, 15], or in systems exhibiting long-wavelength density fluctuations [35, 36]. Conversely, a wide array of out-of-equilibrium systems, such as the fluctuation of red blood cells [37], driven granular gases [16, 33], collections of bristle robots [38], and active colloidal crystals [39], adhere to the GFDR. This observation suggests that, for these systems as well, the Einstein-based effective temperature may hold thermodynamics significance, for example, in contact where energy and particles are exchanged.
@@ Acknowledgements.
The authors thank Prof. Nir Gov of the Weizmann Institute of Science, Prof. Haim Diamant of Tel Aviv University, and Dr. Indrani Chakraborty of the BITS-Pilani Goa campus for their insight and fruitful discussions. YR acknowledges support from the Israel Science Foundation (Grant No. 988/17 and 385/21). ML and YR acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002392).

### v2

@@ Abstract
Our study highlights two key findings in this regime. First, the specific mechanism driving the effective temperature does not influence the relaxation of connected systems toward a steady state. Second, the effective temperature serves a dual purpose: it defines a generalized chemical potential, and together, these two parameters govern the thermalization process in connected systems. In other words, the steady state of the connected systems is reached at maximum entropy, as in thermal equilibrium. Remarkably, this occurs notwithstanding the continuous investment of work and the concomitant entropy production required to sustain a driven, steady state [29].
In this study, our focus was on driven dilute colloidal suspensions. We anticipate that our effective chemical potential may prove inadequate in describing the equilibration of systems in diffusive contact in scenarios where the GFDR is violated. For instance, this could occur in denser driven/active colloidal suspensions [1, 13, 15], or in systems exhibiting long-wavelength density fluctuations [30, 31]. Conversely, a wide array of out-of-equilibrium systems, such as the fluctuations of red blood cells [32], driven granular gases [16, 27], collections of bristle robots [33], and active colloidal crystals [34], adhere to the GFDR. This observation suggests that, for these systems as well, the Einstein-based effective temperature may hold thermodynamic significance, for example, in contact where energy and particles are exchanged.
@@ Acknowledgements.
The authors thank Prof. Nir Gov of the Weizmann Institute of Science, Prof. Haim Diamant of Tel Aviv University, and Dr. Indrani Chakraborty of the BITS-Pilani Goa campus for their insight and fruitful discussions. YR acknowledges support from the Israel Science Foundation (Grant No. 988/17 and 385/21). ML and YR acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101002392).